a number is 3 less than 4 times another number.if their sum is equal to117,then the greater number is










Answer by mananth(16946)

let x & y be the two numbers
x= 4y-3
x-4y=-3............1
x+y=117.............2
Multiply (2) by 4 and add the equation to (1)
4x+4y=468
x-4y=-3
5x=465
x= 465/5
x=93
x= 4y-3
x+3=4y
93+3=4y
96=4y
y=24